OpenStudy (anonymous):

convert the polar form r=3acos(alpha) into Cartesian form, and prove it's a circle

OpenStudy (asnaseer):

\[\begin{align} r&=3a\cos(\alpha)\\ \therefore \cos(\alpha)&=\frac{r}{3a}\\ \text{and }\sin(\alpha)&=\sqrt{1-\cos^2(\alpha)}\\ &=\sqrt{1-\frac{r^2}{9a^2}}\\ &=\frac{\sqrt{9a^2-r^2}}{3a}\\ x&=r\cos(\alpha)\\ &=\frac{r^2}{3a}\\ \therefore r^2&=3ax\\ y&=r\sin(\alpha)\\ &=\frac{r\sqrt{9a^2-r^2}}{3a}\\ \therefore y^2&=\frac{r^2(9a^2-r^2)}{9a^2}\\ &=\frac{3ax(9a^2-3ax)}{9a^2} \end{align}\] can you prove the rest from there?

OpenStudy (anonymous):

thank you ! but can you just instruct me how to prove it ?

OpenStudy (asnaseer):

ok - the final equation can be further simplified as follows...

OpenStudy (asnaseer):

\[\begin{align} y^2&=\frac{9a^2x(3a-x)}{9a^2}\\ &=x(3a-x)\\ &=3ax-x^2\\ \therefore x^2+y^2-3ax&=0\\ \end{align}\]now note that \((x-\frac{3a}{2})^2=x^2-3ax+\frac{9a^2}{4}\), which means we have \(x^2-3ax=(x-\frac{3a}{2})^2-\frac{9a^2}{4}\). so, using this we get...

OpenStudy (asnaseer):

\[\begin{align} x^2+y^2-3ax&=0\\ \therefore (x-\frac{3a}{2})^2-\frac{9a^2}{4}+y^2&=0\\ \therefore (x-\frac{3a}{2})^2+y^2&=\frac{9a^2}{4}&=(\frac{3a}{2})^2 \end{align}\]

OpenStudy (asnaseer):

that is the equation of a circle with origin at \((\frac{3a}{2}, 0)\) and radius \frac{3a}{2}

OpenStudy (asnaseer):

radius = \(\frac{3a}{2}\)

OpenStudy (asnaseer):

do you understand the steps?

OpenStudy (asnaseer):

the general equation of a circle with center at (a, b) and radius r is:\[(x-a)^2+(y-b)^2=r^2\]

OpenStudy (anonymous):

yes i do, thank you very much .. but one thing, how did you know to do it ? because i tried to do \[x=3acos ^{2}\alpha, y=3acos \alpha \sin \alpha\] and then do \[x ^{2}+y ^{2}\] but then i stayed with the alpha variable

OpenStudy (asnaseer):

basically you need to find a way of eliminating both r and alpha. this will come with practice as you do more and more examples.

OpenStudy (asnaseer):

in this case I first got an expression for x interms of just r. then found y in terms of just r. and then looked to see how I could eliminate r from the equations.

OpenStudy (anonymous):

nice, thank you very much !! i definitely need some more practice with it =]

OpenStudy (asnaseer):

one approach is to follow this strategy: 1. you are given r interms of alpha. 2. rearrange this to get an expression for cos/sin alpha in terms of r. 3. you know x = rcos(alpha) - so replace cos(alpha) with the outcome of step 2 4. you know y = rsin(alpha) - so replace sin(alpha) with the outcome of step 2 5. rearrange step 3 to get r in terms of x and substitute this into step 4 and voila!

OpenStudy (anonymous):

great ! thank you very much !!! saved my homework =]

OpenStudy (asnaseer):

yw :-)